OpenStudy (anonymous):

PLEASE HELP WILL MEDAL One estimate of the population of the world on January 1, 2005, is 6,486,915,022. The population is estimated to be increasing at the rate of 1.4% per year. At this rate, what will the population of the world be on January 1, 2024? Round your answer to the nearest whole number. A. 8,212,434,418 B. 8,448,105,986 C. 7,395,083,125 D. 9,086,227,907

OpenStudy (anonymous):

just multiply 6486...*1.014^n where n is diff of 2005 and 2024

OpenStudy (anonymous):

A = P(1 + r/100)^n A = 6,486,915,022*(1 + 1.4/100)^19 A = 8,448,105,986.35075 ~ 8,448,105,986

OpenStudy (anonymous):

Thank you